# Multimodal Unlearnable Examples: Protecting Data against Multimodal Contrastive Learning

### v1

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3680708isbn: 979-8-4007-0686-8/24/10ccs: Security and privacy Privacy protectionsccs: Computing methodologies Computer visionccs: Computing methodologies Machine learning
@@ 1. Introduction
In recent years, there has been a growing interest in multimodal models among researchers in the community (Bengio et al., 2013). Traditional methods (Ma et al., 2024; Li et al., 2024; liang2024object) have primarily focused on analyzing a single modal of data. However, with the rise of multimodal learning, different types of data, such as text, images, and audio, are being combined into a unified framework. One of the most popular approaches for multimodal learning is Multimodal Contrastive Learning (MCL), as demonstrated by models such as CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021). These models are trained with a contrastive loss, which encourages the correlation between pairs of images and captions while also keeping them distinct from unrelated pairs. This approach reduces the need for extensive manual annotation of training data and allows the use of larger datasets that contain millions of examples. MCL has shown promise in various applications, including image classification (Radford et al., 2021), image captioning (Laina et al., 2019; Mokady et al., 2021), image generation (Li et al., 2022b; Patashnik et al., 2021).
@@ 2.2. Poisoning Attacks
Data poisoning attacks aim to disrupt the model training process by processing the training dataset, resulting in a significant increase in test errors for some specific samples during the testing phase. (Shafahi et al., 2018; Chen et al., 2017; Liu et al., 2022). A common type of data poisoning attack is the backdoor attack (Gu et al., 2019; Li et al., 2022a; Liu et al., 2023a; Xun et al., 2024; Gao et al., 2024; Huang et al., 2024b; Liang et al., 2023; Liu et al., 2023b, a; Liang et al., 2024a, b; Zhang et al., 2024; Zhu et al., 2024; Ma et al., 2021, 2022), which typically involves injecting triggers into training samples, leading to misclassification of images containing these trigger patterns during testing. However, it typically only affects samples with trigger, while clean samples remain unaffected and can be correctly classified. Recent work also explores poisoning attacks on multimodal contrastive learning (MCL).  Yang et al. (2023) study the poisoning attacks against multimodal models in both visual and linguistic modalities.  Carlini and Terzis (2021) introduced a framework that effectively poisoned CLIP models with backdoor attacks. In addition, some works aim to train a robust CLIP model against data poisoning and backdoor attacks (Bansal et al., 2023; Yang et al., 2024). Due to the time-consuming of training a high-performance model with a large-scale dataset, we will follow the experimental setting of these works to train a rather simpler model with a small dataset in this paper. When the goal is generalized to the entire test set, the poisoning attack is called an indiscriminate/availability poisoning attack (Yu et al., 2022; He et al., 2022). In the community, this is considered a special case of a poisoning attack, which can also be divided into clean-label and dirty-label scenario.
@@ 2.3. Unlearnable Methods
There are two categories to generate unlearnable examples in image classification. The first category involves model-based attacks, which require a surrogate model to guide the perturbation generation (Shen et al., 2019; Feng et al., 2019). The error minimization (EM)  (Huang et al., 2021b) minimizes the classification error of images on a surrogate classifier and iteratively updates the surrogate model with these perturbed images. Fowl et al. (2021) propose adversarial poisoning, which contains targeted adversarial perturbation (TAP) and untargeted adversarial perturbation (UAP), which uses adversarial examples (Huang et al., 2021a; Gu et al., 2022, 2023; Huang et al., 2024a) as poisoned data to make the model unlearn the features. TAP presents an effective protection for the image dataset. While model-based methods are potent, they are often computationally demanding. Some findings reveal that these approaches can be easily neutralized by adversarial training (AT) (Bai et al., [n. d.]; Jia et al., 2022; Li et al., 2022c; Jia et al., 2024a, b). To address this, several works attempt to leverage robust protection against AT (Fu et al., 2022; Fang et al., 2023). Model-free methods do not require a surrogate model to generate noise for images.  Yu et al. (2022) empirically investigate various unlearnable methods and show that all of them use these spurious features to create a shortcut in the model. They also propose the Linear-sperate Synthetic Perturbation (LSP) in response to this characteristic and show great effectiveness. Autoregressive poisoning (Sandoval-Segura et al., 2022) proposes a generic perturbation generation with each class that can be applied to different datasets, which is a series of dataset-independent perturbations. CUDA (Sadasivan et al., 2023) is generated using controlled class-wise convolutions with filters that are randomly generated via a private key. These previous works focus on image classification using cross-entropy loss, and  He et al. (2022) explore the unlearnable examples for unsupervised contrastive learning and discover that extended EM and TAP methods can still effectively protect against unsupervised learning. However, when extending them to multimodal contrastive learning, their effectiveness diminishes due to the increased modality.
@@ 3.2. Limitation of Existing Works
However, from Table 1, we experimentally found that although EM and UAP can be applied to image-caption pairs, they fail to achieve highly effective protection, especially UAP. We explore the reasons for the decline in the effectiveness of these methods from image classification to multimodal contrastive learning. In image classification, EM and TAP optimize the images with the same label to converge in a feature space, resulting in the model easily capturing these additive noises and learning the correlation with labels, as shown in Fig.2(a). However, in multimodal contrastive learning (MCL), to effectively apply the EM and UAP methods, the direction of the optimized image noise must relate to the features of the caption, causing the image features to become either close to or far away from these features. Nevertheless, the caption features of different pairs may be widely dispersed in the image-caption dataset. Consequently, as illustrated in Fig. 2 (b) and (c), it becomes more challenging for the model to capture the correlation between captions and noise generated by EM and TAP compared to those in classification. In Fig. 2 (c), the learning decision space of UAP is much more complex, so its poor protection is to be expected.
@@ 3.3. Multi-step Error Minimization (MEM)
To this end, we take the EM method as the basic framework and propose adding an additional short text trigger to the caption to minimize contrastive loss, following the setting of an adversarial attack on text tasks (Wallace et al., 2019). Regarding the length of the trigger, longer triggers are more effective, while shorter triggers are more stealthy. Therefore, our method can be conceptualized as a tri-level iterative optimization problem, resembling a multi-step process of EM. Specifically, we sequentially optimize noise $\delta$ and text trigger $t$ to reduce contrastive loss between optimized images $I+\delta$ and optimized text $T\oplus t$, where $\oplus$ denotes the triggers that can be inserted into clean text $T$ at various positions. For simplicity, we choose to add text triggers at the beginning of the text in this paper. As a result, our Multi-step Error Minimization (MEM) method can be formulated as follows:
@@
where $\mathcal{V}$ is the set of all token embeddings in the vocabulary and $\nabla_{\mathbf{e}_{t}}$ is the gradient of loss w.r.t $e_{i}$. We can calculate a set of candidate tokens $e^{\prime}_{j}$ using a dot product with embedding of the vocabulary token and gradients. Finally, we can search for each optimal text trigger with a beam search in a set of candidate tokens. We consider the top-k candidates from Eq. 6 and search front to the last in every position in the trigger and score each beam using the loss on the current batch. We follow  Wallace et al. (2019) and use a small beam size for efficient computation. The pseudocode of our method is shown in Algorithm 1. In Fig. 3, we can see the framework of the generation of multimodal unlearnable examples with our MEM.
@@ 4.1. Experiment Setup
Target models and datasets. Following previous works (Bansal et al., 2023; Yang et al., 2023, 2024), we adopt the open-source implementation of CLIP in the community [^1] https://github.com/mlfoundations/open_clip. For the architecture of the surrogate model, we choose the model with ResNet50 (He et al., 2016) as the image encoder and a Transformer (Vaswani et al., 2017) with certain architecture modifications serving as the text encoder. Our experiment involves three datasets: Flickr8K (Young et al., 2014), Flickr30K (Young et al., 2014), and MS-COCO (Chen et al., 2015). Specifically, Flickr8K and Flickr30K consist of 8000 and 31,000 images, respectively, each accompanied by five captions. We adhere to the protocol established by (Faghri et al., 2017; Karpathy and Fei-Fei, 2015; Zhang et al., 2020), dividing the datasets into training/validation/testing sets with ratios of 6,000/1,000/1,000 and 29,000/1,000/1,000 for Flickr8K and Flickr30K, respectively. MS-COCO comprises 123,287 images, each annotated with five descriptions. As per (Karpathy and Fei-Fei, 2015), MS-COCO underwent division into 82,783 training images, 5,000 validation images, and 5,000 test images. Despite their smaller size compared to the large-scale datasets used to train the original CLIP model, these datasets remain suitable for storage and computational resources and have found widespread usage in MCL studies (Yang et al., 2023, 2024).
@@
Evaluation Metrics. We evaluate the impact of data protection by examining the decline in performance in image and text retrieval tasks, drawing from previous work on attacks in MCL (Yang et al., 2024, 2023). Two metrics are employed to illustrate the retrieval results. Hit@10 measures the proportion of all target images/texts that appear within the top 10 of the list of rank. Higher Hit@10 values indicate that many text/image samples successfully retrieving target images/texts early, reflecting a better rank list. Medr refers to the median position of all target images / texts in the list of test images/texts. Lower Rank values signify earlier access to target images, indicative of a superior rank list. The performance of unlearnable multimodal examples is assessed using Hit@10 and Medr for image retrieval (Image $\rightarrow$ Text) and text retrieval (Text $\rightarrow$ Image) across all testing images. Lower Hit@10 and higher Medr-Rank values signify more effective protection of the data.
@@ 4.2. Effectiveness and Transferability
Fig. 4 illustrates the descending curves of the training loss trained on unlearnable examples generated by different methods and the retrieval Medr on the clean test set. From (a), we observe that although EM enables the loss to fall faster compared to normal training, our method, MEM-3 and MEM-5, have a smaller loss on the first epoch, suggesting that the model can quickly learn the shortcuts. The UAP hardly speeds up the loss decline significantly, so it is understandably inefficient. From (b), we find that all models are trained with Medr decreasing compared to random guessing, but the model trained on unlearnable examples generated by MEM stops learning the fastest, reaches the worst retrieval result, and does not learn better further as the epoch increases. The above observations are consistent with the results in Table 1.
We assume that data protection is a completely black-box setting, wherein the protector lacks knowledge of the hacker model’s architecture. Thus, we evaluate the performance of our MEM generated on the ResNet50 surrogate model on different hacker models, including ResNet101, and ViT-B/32. The results are presented in Table 2. We find that these examples can be successfully transferred across different models and can degrade the performance of the CLIP model. Although they generated a surrogate model with a structure different from that of the hacker model can also exhibit the same protective effect, this observation aligns with the one of unlearnable examples in image classification. This is likely because the feature preferences learned by the models are similar.
4.3. Analysis
Attention of the models. Fig. 5 presents the heatmaps of images and text, illustrating the attention of models trained on clean and unlearnable examples. The Grad-CAM (Selvaraju et al., 2017) is utilized to visualize the model attention for images, while the Integrated Gradients (Sundararajan et al., 2017) is employed to visualize the attention to the text. Lighter colors represent higher attention from the model. Remarkably, for the image, models in (a), (b), and (c) all focus on the central region, which correlates with the captions. In contrast, model (d), trained on samples generated by MEM-3, fails to accurately recognize the clean image due to only learning noise features. Similarly in the text, models in (a), (b), and (c) all focus on the keywords of ’glass’, while the model in (d) put the attention on the first three words, probably because MEM-3 always optimizes the noise and the first three text triggers to create shortcuts. These visualizations report that the EM and UAP are ineffective enough in protecting the multimodal data, while MEM-3 has an obvious effectiveness.
(a) Clean
(b) MEM-3
Extension with Semantic Triggers. In the previous methodology, we propose to use the gradient by the Hotflip method to select embeddings in the vocabulary list with large inner products to serve as candidate words for replacement. However, this approach may result in the generation of candidate words lacking semantic information, potentially impacting social reading or being detected and removed by hackers. Therefore, here we leverage the pre-training parameters of the BERT model to generate some semantically relevant substitutions as vocabulary lists. Subsequently, we embed these substitutions to obtain $\mathcal{V}_{\text{bert}}$, and then compute the inner product with the gradient and token embeddings of the substitutions, which yields replaced words with semantics. Table 3 displays examples of text triggers generated by our MEM-3 method under different datasets, including those generated with and without the BERT method. Text triggers generated with BERT exhibit greater naturalness and semantic coherence compared to those generated without BERT. Additionally, we show the Medr results of these triggers in text-image retrieval. It is observed that triggers generated without BERT have better protection performance, which is attributed to their larger search space, while triggers generated with BERT are restricted to a vocabulary space $\mathcal{V}_{\text{bert}}$, resulting in inferior protection efficacy.

### v2

@@ Abstract.
doi: 10.1145/3664647.3680708journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28–November 1, 2024; Melbourne, VIC, Australia.booktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28–November 1, 2024, Melbourne, VIC, Australiaisbn: 979-8-4007-0686-8/24/10ccs: Security and privacy Privacy protectionsccs: Computing methodologies Computer visionccs: Computing methodologies Machine learning
@@ 1. Introduction
In recent years, there has been a growing interest in multimodal models among researchers in the community (Bengio et al., 2013). Traditional methods (Ma et al., 2024; Li et al., 2024; Liang et al., 2024d) have primarily focused on analyzing a single modal of data. However, with the rise of multimodal learning, different types of data, such as text, images, and audio, are being combined into a unified framework. One of the most popular approaches for multimodal learning is Multimodal Contrastive Learning (MCL), as demonstrated by models such as CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021). These models are trained with a contrastive loss, which encourages the correlation between pairs of images and captions while also keeping them distinct from unrelated pairs. This approach reduces the need for extensive manual annotation of training data and allows the use of larger datasets that contain millions of examples. MCL has shown promise in various applications, including image classification (Radford et al., 2021), image captioning (Laina et al., 2019; Mokady et al., 2021), image generation (Li et al., 2022a; Patashnik et al., 2021).
@@ 2.2. Poisoning Attacks
Data poisoning attacks aim to disrupt the model training process by processing the training dataset, resulting in a significant increase in test errors for some specific samples during the testing phase. (Shafahi et al., 2018; Chen et al., 2017; Liu et al., 2022). A common type of data poisoning attack is the backdoor attack (Gu et al., 2019; Liu et al., 2024; Xun et al., 2024; Gao et al., 2024; Huang et al., 2024b; Liang et al., 2023; Liu et al., 2023; Liang et al., 2024a, b; Zhang et al., 2024; Zhu et al., 2024; Ma et al., 2021, 2022), which typically involves injecting triggers into training samples, leading to misclassification of images containing these trigger patterns during testing (Gao et al., 2023; Lan et al., 2024). However, it typically only affects samples with trigger, while clean samples remain unaffected and can be correctly classified. Recent work also explores poisoning attacks on multimodal contrastive learning (MCL).  Yang et al. (2023) study the poisoning attacks against multimodal models in both visual and linguistic modalities.  Carlini and Terzis (2021) introduced a framework that effectively poisoned CLIP models with backdoor attacks. In addition, some works aim to train a robust CLIP model against data poisoning and backdoor attacks (Bansal et al., 2023; Yang et al., 2024). Due to the time-consuming of training a high-performance model with a large-scale dataset, we will follow the experimental setting of these works to train a rather simpler model with a small dataset in this paper. When the goal is generalized to the entire test set, the poisoning attack is called an indiscriminate/availability poisoning attack (Yu et al., 2022; He et al., 2022). In the community, this is considered a special case of a poisoning attack, which can also be divided into clean-label and dirty-label scenario.
@@ 2.3. Unlearnable Methods
There are two categories to generate unlearnable examples in image classification. The first category involves model-based attacks, which require a surrogate model to guide the perturbation generation (Shen et al., 2019; Feng et al., 2019). The error minimization (EM)  (Huang et al., 2021b) minimizes the classification error of images on a surrogate classifier and iteratively updates the surrogate model with these perturbed images.  Fowl et al. (2021) propose adversarial poisoning, which contains targeted adversarial perturbation (TAP) and untargeted adversarial perturbation (UAP), which uses adversarial examples (Huang et al., 2021a; Gu et al., 2021, 2022b, 2023; Huang et al., 2024a; Gu et al., 2022a) as poisoned data to make the model unlearn the features. TAP presents an effective protection for the image dataset. While model-based methods are potent, they are often computationally demanding. Some findings reveal that these approaches can be easily neutralized by adversarial training (AT) (Bai et al., [n. d.]; Jia et al., 2022; Li et al., 2022b; Jia et al., 2024a, b). To address this, several works attempt to leverage robust protection against AT (Fu et al., 2022; Fang et al., 2023). Model-free methods do not require a surrogate model to generate noise for images.  Yu et al. (2022) empirically investigate various unlearnable methods and show that all of them use these spurious features to create a shortcut in the model. They also propose the Linear-sperate Synthetic Perturbation (LSP) in response to this characteristic and show great effectiveness. Autoregressive poisoning (Sandoval-Segura et al., 2022) proposes a generic perturbation generation with each class that can be applied to different datasets, which is a series of dataset-independent perturbations. CUDA (Sadasivan et al., 2023) is generated using controlled class-wise convolutions with filters that are randomly generated via a private key. These previous works focus on image classification using cross-entropy loss, and  He et al. (2022) explore the unlearnable examples for unsupervised contrastive learning and discover that extended EM and TAP methods can still effectively protect against unsupervised learning. However, when extending them to multimodal contrastive learning, their effectiveness diminishes due to the increased modality.
@@ 3.2. Limitation of Existing Works
However, from Table 1, we experimentally found that although EM and UAP can be applied to image-caption pairs, they fail to achieve highly effective protection, especially UAP. We explore the reasons for the decline in the effectiveness of these methods from image classification to multimodal contrastive learning. In image classification, EM and UAP optimize the images with the same label to converge in a feature space, resulting in the model easily capturing these additive noises and learning the correlation with labels, as shown in Fig.2(a). However, in multimodal contrastive learning (MCL), to effectively apply the EM and UAP methods, the direction of the optimized image noise must relate to the features of the caption, causing the image features to become either close to or far away from these features. Nevertheless, the caption features of different pairs may be widely dispersed in the image-caption dataset. Consequently, as illustrated in Fig. 2 (b) and (c), it becomes more challenging for the model to capture the correlation between captions and noise generated by EM and UAP compared to those in classification. In Fig. 2 (c), the learning decision space of UAP is much more complex, so its poor protection is to be expected.
@@ 3.3. Multi-step Error Minimization (MEM)
To this end, we take the EM method as the basic framework and propose adding an additional short text trigger to the caption to minimize contrastive loss, following the setting of an adversarial attack on text tasks (Wallace et al., 2019). Regarding the length of the trigger, longer triggers are more effective, while shorter triggers are more stealthy. Therefore, our method can be conceptualized as a tri-level iterative optimization problem, resembling a multi-step process of EM. Specifically, we sequentially optimize noise $\delta$ and text trigger $t$ to reduce contrastive loss between optimized images $I+\delta$ and optimized text $T\oplus t$, where $\oplus$ denotes the triggers that can be inserted into clean text $T$ at various positions. For simplicity, we choose to add text triggers at the beginning of the text. As a result, our Multi-step Error Minimization (MEM) can be formulated as follows:
@@
where $\mathcal{V}$ is the set of all token embeddings in the vocabulary and $\nabla_{\mathbf{e}_{t}}$ is the gradient of loss w.r.t $e_{i}$. We can calculate a set of candidate tokens $e^{\prime}_{j}$ using a dot product with embedding of the vocabulary token and gradients. Finally, we can search for each optimal text trigger with a beam search in a set of candidate tokens. We consider the top-k candidates from Eq. 6 and search front to the last in every position in the trigger and score each beam using the loss on the current batch. We follow  Wallace et al. (2019) and use a small beam size for efficiency. In Fig. 3, we can see the framework of the generation of multimodal unlearnable examples with our MEM.
@@ 4.1. Experiment Setup
Target models and datasets. Following previous works (Bansal et al., 2023; Yang et al., 2023, 2024), we adopt the open-source implementation of CLIP in the community [^1] https://github.com/mlfoundations/open_clip. For the architecture of the surrogate model, we choose the model with ResNet50 (He et al., 2016) as the image encoder and a Transformer (Vaswani et al., 2017) with certain architecture modifications serving as the text encoder. Our experiment involves three datasets: Flickr8K (Young et al., 2014), Flickr30K (Young et al., 2014), and MS-COCO (Chen et al., 2015). Specifically, Flickr8K and Flickr30K consist of 8000 and 31,000 images, respectively, each accompanied by five captions. We adhere to the protocol established by (Faghri et al., 2017; Karpathy and Fei-Fei, 2015; Zhang et al., 2020), dividing the datasets into training/validation/testing sets with ratios of 6,000/1,000/1,000 and 29,000/1,000/1,000 for Flickr8K and Flickr30K, respectively. MS-COCO comprises 123,287 images, each annotated with five descriptions. In  (Karpathy and Fei-Fei, 2015), MS-COCO underwent division into 82,783 training images, 5,000 validation images, and 5,000 test images. Despite their smaller size compared to the large-scale datasets used to train the original CLIP model, these datasets remain suitable for storage and computational resources and have found widespread usage in MCL studies (Yang et al., 2023, 2024).
@@
Evaluation Metrics. We evaluate the impact of data protection by examining the decline in performance in image and text retrieval tasks, drawing from previous work on attacks in MCL (Yang et al., 2024, 2023). Two metrics are employed to illustrate. Hit@10 measures the proportion of all target images/texts that appear within the top 10 of the list of rank. Higher Hit@10 values indicate that many text/image samples successfully retrieving target images/texts early, reflecting a better rank list. Medr refers to the median position of all target images / texts in the list of test images/texts. Lower Rank values signify earlier access to target images, indicative of a superior rank list. The performance of unlearnable multimodal examples is assessed using Hit@10 and Medr for image retrieval (Image $\rightarrow$ Text) and text retrieval (Text $\rightarrow$ Image) across all testing images. Lower Hit@10 and higher Medr signify more effective protection of the data.
@@ 4.2. Effectiveness and Transferability
Fig. 4 illustrates the descending curves of the training loss trained on unlearnable examples generated by different methods and the retrieval Medr on the clean test set. From (a), we observe that although EM enables the loss to fall faster compared to normal training, our method, MEM-3 and MEM-5, have a smaller loss on the first epoch, suggesting that the model can quickly learn the shortcuts. The UAP hardly speeds up the loss decline significantly, so it is understandably inefficient. From (b), we find that all models are trained with Medr decreasing compared to random guessing, but the model poisoned by MEM stops learning the fastest, reaches the worst result, and does not learn further as the epoch increases. The above observations are consistent with the results in Tab. 1.
We assume that data protection is a completely black-box setting, wherein the protector lacks knowledge of the hacker model’s architecture. Thus, we evaluate the performance of our MEM generated on the ResNet50 surrogate model on different hacker models, including ResNet101, and ViT-B/32. The results are presented in Tab. 2. We find that these examples can be successfully transferred across different models and can degrade the performance of the CLIP model. Although they generated a surrogate model with a structure different from that of the hacker model can also exhibit the same protective effect, this observation aligns with the one of unlearnable examples in image classification. This is likely because the feature preferences learned by the models are similar.
4.3. Analysis
Attention of the models. Fig. 5 presents the heatmaps of images and text, illustrating the attention of models trained on clean and unlearnable examples. The Grad-CAM (Selvaraju et al., 2017) is utilized to visualize the model attention for images, while the Integrated Gradients (Sundararajan et al., 2017) is employed to visualize the attention to the text. Lighter colors represent higher attention from the model. Remarkably, for the image, models in (1), (2), and (3) all focus on the central region, which correlates with the captions. In contrast, model (4), trained on samples generated by MEM-3, fails to accurately recognize the clean image due to only learning noise features. Similarly in the text, models in (1), (2), and (3) all focus on the keywords of ’glass’, while the model in (d) put the attention on the first three words, probably because MEM-3 always optimizes the noise and the first three text triggers to create shortcuts. These visualizations report that the EM and UAP are ineffective enough in protecting the multimodal data, while MEM-3 has an obvious effectiveness.
(a) Clean
(b) MEM-3
@@ 4.3. Analysis
Extension with Semantic Triggers. In the previous parts, we propose to use the gradient by the Hotflip method to select embeddings in the vocabulary list with large inner products to serve as candidate words for replacement. However, this approach may result in the generation of candidate words lacking semantic information, potentially impacting social reading or being removed by hackers. Here, we leverage the pre-training parameters of the BERT model to generate some semantically relevant substitutions as vocabulary lists. Subsequently, we embed these substitutions to obtain $\mathcal{V}_{\text{bert}}$, and then compute the inner product with the gradient and token embeddings of the substitutions, which yields replaced words with semantics. Table 3 displays examples of text triggers generated by our MEM-3 method under different datasets, including those generated with and without the BERT method. Text triggers generated with BERT exhibit greater naturalness and semantic coherence compared to those generated without BERT. Additionally, we show the Medr results of these triggers in text-image retrieval. It is observed that triggers generated without BERT have better protection performance, which is attributed to their larger search space, while triggers generated with BERT are restricted to a vocabulary space $\mathcal{V}_{\text{bert}}$, resulting in inferior protection efficacy.